# UMArm: Untethered, Modular, Wearable, Soft Pneumatic Arm

### v1

@@ III-A Forward Kinematics of the UMArm
The arm is modeled as a 12-revolute joint robot. In this work, we employ the standard product of exponential method [40]. For each U-joint, We define the two revolute joints as illustrated in Fig. 2B. The rotation axes of each of the 12 revolute joints can be expressed as $\overrightarrow{\omega}=[\omega_{0},...,\omega_{11}]^{\intercal},\omega_{i}\in\mathbb{R}^{3}$, and we choose $q_{i}$ as the center of the U-joint, where two revolute axes intersect: $\overrightarrow{q}=[q_{0},...,q_{11}]^{\intercal},q_{i}\in\mathbb{R}^{3}$. The rotation angle variable for each revolute joint is $\overrightarrow{\theta}=[\theta_{0},...,\theta_{11}]^{\intercal},\theta_{i}\in\mathbb{R}$. The 3 segments of the arm have identical structure, therefore, we can first compute the forward kinematics for each segment consists of 4 joints using $\overrightarrow{\omega}$, $\overrightarrow{q}$, and $\overrightarrow{\theta}$:
@@
$g_{st}^{robot}(\theta)=g_{st}^{seg3}(\theta_{8-11})H_{s_{2}}^{s_{3}}g_{st}^{seg2}(\theta_{4-7})H_{s_{1}}^{s_{2}}g_{st}^{seg1}(\theta_{0-3})$ (2)
$g_{st}^{robot}(\theta)$ will yield the tool-frame position and orientation in $SE(3)$ format expressed in the robot coordinate frame. Robot specific dimensional parameters and more detail regarding the calculation process can be found in the Supplementary Materials. In the actual implementation, we use Wolfram Mathematica to pre-compute one analytic expression of the products of exponential for all three segments. It is a function of segment dimension parameters and segment joint variables. We exported this function to Python for robot control. Kinematic code can be found in the following GitHub repository (link)
@@ III-B Inverse Kinematics of the UMArm
The inverse kinematics of UMArm is achieved through using the iterative inverse Jacobian method [40]. Each revolute joint on the arm has a ±$\pm 15\degree$ range of motion. In this work, the inverse kinematics problem will be solved for the position of the end-effector, while we assume that the orientation will be achieved with extra DoFs on the end-effector. “Vanilla” Jacobian inverse kinematics cannot limit the joint angle or smartly distribute motion among the 12 DoF. As a result, the IK solver will output solutions that are not feasible for our physical system. Modifications need to be implemented to the Jacobian IK solver. UMArm has many redundancies granted by the large number of independently actuated muscles. This will make the null space of the manipulator Jacobian non-empty, which effectively means that the robot can change its posture in specific ways while still maintaining the end-effector position. We can use this feature to our advantage to redistribute motion reasonably among the joints. We have implemented a method that combines the gradient projection method [41] and the weighted least norm method [42]. These modifications will encourage each joint to stay within the range of motion limit. The detailed Jacobian derivation and the gradient projection method can be found in Supplementary Materials
Figure 3: Closed loop controller diagram of the joint PID controller and the actuator internal pressure on-off controller.
@@ III-C Joint Angle Controller
Each revolute joint is designed to be controlled by two VEMAs as described in Fig. 2C. Since the actuators are antagonistically installed, the stiffness of each joint can be tuned by simultaneously actuating the two actuators. The arm user can set the target antagonistic stiffness for each joint as $p_{A}$, where $p_{A}$ will be the minimum pressure for the antagonistic pair of actuators of that joint. We use a standard PID controller for each revolute joint to control the joint angle in a closed-loop manner. As shown in the control diagram in Fig. 3, the PID controller acquires the actual joint angle feedback from the motion capture system. According to the error input, the PID controller will decide the ratio $p_{ratio}=p1/p2$, where $p_{1}$ and $p_{2}$ are the pressures in antagonistic VEMAs. Then, depending on the rotation direction, either $p_{1}$ or $p_{2}$ will be set to $p_{A}$, and the other one will be calculated using $p_{ratio}$.
The antagonistic pressure $p_{A}$ of all 12 revolute joints can be set independently, which opens up new possibilities for the control of the arm. However, applying antagonistic stiffness to the joint will decrease its range of motion. Therefore, we have a heuristic that will automatically lower the antagonistic pressure when the joint is trying to reach a larger angle. Detailed controller tuning can be found in Supplementary Materials.
Figure 4: Schematic of the experimental setup. The actuators (from our previous work [25]) are connected to a bus line that receives voltage from a battery pack, pneumatic pressure from a compressed air tank, and command signals wirelessly via Bluetooth. The arm posture is captured by a 12-camera motion capture system. A laptop is running a Python application that processes the arm posture stream and formulates arm control commands.
@@ IV-A Experiment Setup
In the characterization experiments, the arm’s base is mounted to the ceiling rail, with the arm naturally hanging down as the rest configuration. The arm and its peripheral system are connected as shown in Figure 4. The arm operates using a laboratory compressed air outlet and a 12V DC source. The two-line RS485 serial communication interface connects to an RS485-to-USB adapter, then plugs into the PC. A 12-camera motion capture system is used to track the deformation of the arm. Motion capture markers are installed on cantilever arms on the printed link frame shown in Fig. 2C. Four markers on one link frame define one rigid body. In each rigid body, the four markers form a quadrilateral, and the intersection point of its diagonals aligns with the center of the U-joint. The 3-segment arm has six rigid bodies in total, and the rigid bodies’ positions and orientations can be acquired from the motion capture system. By using the rigid body position and orientation, we can compute each joint angle defined in Figure 2A. A multi-process Python application runs on the laptop PC. One process receives the data stream from the motion capture server, and for each frame, calculates the joint angle of the robot. Another process runs the 12-joint PID controller and the serial interface that sends target pressures to the actuators. A third process is dedicated to the computationally heavy kinematics functions, and the last process runs an interface that receives user commands and visualizes the robot. In the characterization section, a 10 cm rod with 4 motion capture markers (shown in Fig.2A) is installed as the end-effector.
@@ IV-B Step Response of the Arm
The joint angle controller’s performance is characterized by a step response experiment. At $t=0s$, the joint angle controller is enabled, and all 12 joints are set to zero. At $t=6s,12s,18s$, all joints are set to $-0.1,0,0.1\ (rad)$ respectively. For better figure readability, we compute the forward kinematics using the target joint angles and the actual joint angles, and plot the end-effector position vs time in figure 5. Another figure with 12 joint angle subplots is available in Supplementary Materials
Figure 5: Step response plot for the arm’s joint angle controller. Every 6s, the target joint angle is switched between 0,0.1,-0.1 rad. For more intuitive reading, target and actual joint angles are converted to end-effector spatial coordinates using forward kinematics. The experiment is repeated three times for different antagonistic stiffening and PID tuning profiles.
Antagonistic/PID 5% Settling Tracking RMSE
High/Aggressive 1.32s 0.36°
Low/Aggressive Does not settle 0.81°
Low/Conservative Oscillate 1.09°
Figure 6: PID controller performance of three different antagonistic stiffening and PID tuning profiles.
To visualize the benefit of having antagonistic stiffening at each joint, we performed the step response experiment with three different stiffnesses and controller tunings. We first used a low antagonistic pressure $p_{A}$=14 kPa (2 psi) and fine-tuned the PID gains with our best effort to minimize the tracking error while avoiding excessive oscillation. We define this PID gain profile as “Conservative”, and the tracking result is plotted in Fig. 5 in the medium blue line. The system cannot settle within 5% in the given timeframe, and the average joint angle tracking RMSE for the whole experiment is 1.09°.
Next, we increased the antagonistic pressure of the first(base) segment to 83 kPa (12 psi), the second segment to 8 psi, and the third segment to 28 kPa (4 psi). The antagonistic pressure is in a decreasing order because the segments closer to the base bear a heavier load. Using this stiffness profile, we can significantly increase the PID gains without introducing oscillations. The newly tuned PID gain profile is defined as “Aggressive”, and the tracking result is plotted in a dark blue line. The 5% settling time is 1.32s, and the average joint angle tracking RMSE for the whole experiment is 0.36°, illustrating a significant performance gain by having antagonistic stiffness.
To demonstrate the oscillatory behavior when larger PID gains are applied to a low structural stiffness system, we use the “Aggressive” tuning profile to the $p_{A}=14\ kPa(2\ psi)$ low antagonistic setting. The oscillatory behavior is illustrated by the light blue line of Fig. 5.
Figure 7: A: Way-point traversing experiment. The arm is given 8 way-points. Using the inverse kinematic solver, the arm will place its end-effector at the way-points. The time interval between each way-point is 5 seconds. B: Photos of UMArm performing way-point tracking. The way-points are marked with red dots. Each photo shows the arm moving its end-effector to two of the eight way-points. White dashed arrows indicate the traversal order.
At “High” antagonistic structural stiffness and “Aggressive” PID tuning, UMArm can achieve a steady-state joint angle RMSE of 0.033 degrees (excluding the error for the first 3s after a new target is given and assuming motion capture data as ground truth for the joint controller).
@@
Qualitatively speaking, a joint with higher antagonistic stiffness will act like a spring with a large spring constant. Upon disturbances, its response will have a lower magnitude. Also, a McKibben actuator at higher pressure can dissipate energy faster due to the larger internal friction between the fibers. Therefore, the joint can structurally (without control) lower the oscillation magnitude and increase the damping factor. As a result, the arm can quickly stabilize when a new joint angle target is given.
IV-C Precision, Payload Capacity, and Untethered Operation
The UMArm’s inverse kinematics controller is characterized by a way-point traversal experiment. The arm is tasked with traversing eight way-points in a specific order, as shown in Figure 7B. Every five seconds, the robot will switch to the next way-point. The inverse kinematics controller is as described in section II. It runs in real time and iteratively updates the target robot configuration. Figure 7A shows the target and the actual position of the end-effector. The inverse kinematics algorithm converges almost instantly, so the tracking error mostly comes from the joint controller. Using the last two seconds from the five-second interval for each way-point, we can compute the average RMSE to be 2.6mm.
Payload capacity is characterized by attaching weights to the tip of the arm and actuating the arm to its maximum bending position. As shown in Figure 8D, we tested payload ranging from 0 to 3 Kg. The range of motion of the arm will be affected by the heavier payload.
@@ IV-C Precision, Payload Capacity, and Untethered Operation
The arm can operate untethered using a portable pressure source. We demonstrate this functionality using an off-the-shelf compressed air tank that holds 1.11L (68ci) at 31MPa (4500 psi) (Fig 10A). It was originally built for airsoft players, so it is designed to be portable and lightweight. We characterize the untethered performance by letting the robot traverse the eight way-points in the previous experiment with two 2-second intervals. This experiment shows that the compressed air tank we use could supply the robot to traverse 700 way points, or run continuously for 11min 45s before it starts to lose tracking accuracy due to the depletion of the tank.
@@ V UMArm Demonstrations
V-A Safe and Dexterous Arm
Figure 8: A: An 2-DoF end-effector module consists of 2 encoder motors and a gripper is installed on the arm. Using the arm’s current configuration, the controller will calculate encoder motor angles so that the motors can maintain the spatial orientation of the gripper. The gripper is driven by a mini servo. B: With the installed gripper end-effector, the arm can stack the three objects together. C: The arm is controlled to be stiffened in one direction and can easily pierce through the paper. At the same time, it is controlled to be compliant in another direction for safer interaction with humans. D: Payloads are connected to the tip of the UMArm while it actuates one side of its muscle to max pressure. The final configurations for each payload level are overlaid on the figure.
@@ V-A Safe and Dexterous Arm
Soft robotic arms are structurally compliant, which is crucial for safe interaction with humans. However, for a robotic arm to be precise and fast, it needs to have control over its structural stiffness. Otherwise, we might see oscillatory behavior or buckling under heavy load. UMArm can control the structural compliance on all 12 joints on the arm. This has granted the UMArm great flexibility to have safety and performance simultaneously. Figure 8C demonstrates UMArm’s directional compliance control capability. Two similarly sharp objects (scissors and a screwdriver) are installed perpendicularly on the end effector. Then, we selectively stiffen some of the joints and relax the other joints to make UMArm structurally stiff along the scissors direction while soft along the screwdriver direction. This directional difference in stiffness can be demonstrated by a human operator holding a piece of paper and pushing it against the two sharp objects. Kitchen wax paper is used on the more rigid direction, while a napkin is used on the soft direction. As shown in Figure 8C, UMArm can poke through the strong wax paper with minimal displacement of the arm, while it cannot poke through a soft napkin along the soft direction.

### v2

@@ III-A Forward Kinematics of the UMArm
The arm is modeled as a 12-revolute joint robot. In this work, we employ the standard product of exponential method [murray2017mathematical]. For each U-joint, We define the two revolute joints as illustrated in Fig. 2B. The rotation axes of each of the 12 revolute joints can be expressed as $\overrightarrow{\omega}=[\omega_{0},...,\omega_{11}]^{\intercal},\omega_{i}\in\mathbb{R}^{3}$, and we choose $q_{i}$ as the center of the U-joint, where two revolute axes intersect: $\overrightarrow{q}=[q_{0},...,q_{11}]^{\intercal},q_{i}\in\mathbb{R}^{3}$. The rotation angle variable for each revolute joint is $\overrightarrow{\theta}=[\theta_{0},...,\theta_{11}]^{\intercal},\theta_{i}\in\mathbb{R}$. The 3 segments of the arm have identical structure, therefore, we can first compute the forward kinematics for each segment consists of 4 joints using $\overrightarrow{\omega}$, $\overrightarrow{q}$, and $\overrightarrow{\theta}$:
@@
$g_{st}^{robot}(\theta)=g_{st}^{seg1}(\theta_{0-3})H_{s_{1}}^{s_{2}}g_{st}^{seg2}(\theta_{4-7})H_{s_{2}}^{s_{3}}g_{st}^{seg3}(\theta_{8-11})$ (2)
$g_{st}^{robot}(\theta)$ will yield the tool-frame position and orientation in $SE(3)$ format expressed in the robot coordinate frame. Robot-specific dimensional parameters and more details regarding the calculation process can be found in the Supplementary Materials. In the actual implementation, we use Wolfram Mathematica to pre-compute one analytic expression of the products of exponentials for all three segments. It is a function of segment dimension parameters and segment joint variables.
@@ III-B Inverse Kinematics of the UMArm
The inverse kinematics of UMArm is achieved through using the iterative inverse Jacobian method [murray2017mathematical]. Each revolute joint on the arm has a $\pm 0.3~rad~(17.2\degree)$ range of motion. In this work, the inverse kinematics problem will be solved for the position of the end-effector, while we assume that the orientation will be achieved with extra DoFs on the end-effector. “Vanilla” Jacobian inverse kinematics cannot limit the joint angle or smartly distribute motion among the 12 DoF. As a result, the IK solver will output solutions that are not feasible for our physical system. Modifications need to be implemented to the Jacobian IK solver. UMArm has many redundancies granted by the large number of independently actuated muscles. This makes the null space of the manipulator Jacobian non-empty, which effectively means that the robot can change its posture in specific ways while still maintaining the end-effector position. We can use this feature to our advantage to redistribute motion reasonably among the joints. We have implemented a method that combines the gradient projection method [liegeois1977automatic] and the weighted least norm method [chan1995weighted]. These modifications will encourage each joint to stay within the range of motion limit. The detailed Jacobian derivation and the gradient projection method can be found in the Supplementary Materials.
Figure 3: A: Closed loop controller diagram of the joint PID controller and the actuator internal pressure on-off controller. B: Mapping from the PID controller output to the pressure inside the two antagonistic pair actuators.
@@ III-C Joint Angle Controller
Each revolute joint is designed to be controlled by two VEMAs as described in Fig. 2C. We use a standard PID controller for each revolute joint to control the joint angle in a closed-loop manner. As shown in the control diagram in Fig. 3A, the PID controller acquires the actual joint angle feedback from the motion capture system. According to the error input and the target joint angle setpoint, the PID controller outputs a value between -1 and 1. Since two actuators control one joint, we need to convert the PID output into actuator pressures.
Figure 3B shows the method of mapping PID output onto actuator pressures. The two curves represent the pressure inside the two antagonistic pair actuators. At zero PID output, both actuators start with the initial antagonistic pressure $p_{antg}$. As the PID output change, depending on the direction, one actuator pressure will increase toward the maximum actuator pressure $p_{max}$, while the other actuator will decrease toward the minimun actuator pressure $p_{min}$. The change rate $k_{p}$ is set to a constant value.
$\begin{cases}p_{\text{A}}=p_{\text{antg}}+k_{\text{p}}\cdot PID_{\text{out}}\\ p_{\text{B}}=p_{\text{antg}}-k_{\text{p}}\cdot PID_{\text{out}}\\ p_{\text{A}}=min(p_{\text{max}},p_{\text{A}})\\ p_{\text{B}}=max(p_{\text{min}},p_{\text{B}})\end{cases}$ (3)
The stiffness of each joint can be tuned by simultaneously (antagonistically) increasing/decreasing the pressure inside the two actuators. To do that, the arm user can set the initial antagonistic stiffness for each joint $p_{antg}$. The mapping shown above is a simple and balanced method for converting PID output to actuator pressures. It is also possible to generate different mapping schemes that fit the need for different applications.
Figure 4: Schematic of the experimental setup. The actuators (from our previous work [zuo2024embedded]) are connected to a bus line that receives voltage from a battery pack, pneumatic pressure from a compressed air tank, and command signals wirelessly via Bluetooth. The arm posture is captured by a 12-camera motion capture system. A laptop is running a Python application that processes the arm posture stream and formulates arm control commands.
@@ IV-A Experiment Setup
In the characterization experiments, the arm’s base is mounted to the ceiling rail, with the arm naturally hanging down as the rest configuration. The arm and its peripheral system are connected as shown in Figure 4. The arm operates using a laboratory compressed air outlet and a 12V DC source. The two-line RS485 serial communication interface connects to an RS485-to-USB adapter, then plugs into the PC. A 12-camera motion capture system (OptiTrack Flex 13) is used to track the deformation of the arm. Motion capture markers are installed on cantilever arms on the printed link frame shown in Fig. 2C. Four markers on one link frame define one rigid body. In each rigid body, the four markers form a quadrilateral, and the intersection point of its diagonals aligns with the center of the U-joint. The 3-segment arm has six rigid bodies in total, and the rigid bodies’ positions and orientations can be acquired from the motion capture system. By using the rigid body position and orientation, we can compute each joint angle defined in Figure 2A. A multi-process Python application runs on the laptop PC. One process receives the data stream from the motion capture server, and for each frame, calculates the joint angle of the robot. Another process runs the 12-joint PID controller and the serial interface that sends target pressures to the actuators. A third process is dedicated to the computationally heavy kinematics functions, and the last process runs an interface that receives user commands and visualizes the robot. In the characterization section, a 0.1m rod with 4 motion capture markers (shown in Fig.2A) is installed as the end-effector.
@@ IV-B Step Response of the Arm
Figure 5: Position control step response of the UMArm. At $t=0$, the arm starts from its resting position and is commanded to reach a target position using the inverse kinematics controller. Four load conditions are tested. For each load, the effects of antagonistic pressure and controller gain are evaluated. The 5% settling times are summarized in the accompanying table, and for each antagonistic pressure setting, the best-performing trial is plotted. Line colors correspond to those in the table.
The performance of the joint angle controller and the inverse kinematics position controller is evaluated through a step response experiment. At $t=0$ s, the inverse kinematics controller is activated (which also utilizes the joint angle controller), and the target position is set to the position at the initial configuration with all joints at zero. Subsequently, a new target position located 0.25 m from the initial position is commanded, and the UMArm attempts to reach this position with its end-effector.
Figure 5 presents the results of the step response experiment under four loading conditions: 0 kg, 0.5 kg, 1 kg, and 2 kg. To evaluate the effect of antagonistic stiffening, three levels of structural stiffness were tested for each load by varying the antagonistic pressure ($p_{\text{antg}}$) to 110 kPa, 55 kPa, and 14 kPa. In each case, the chosen $p_{antg}$ was applied uniformly across all 12 pairs of antagonistic actuators.
[…]
@@
Qualitatively speaking, a joint with higher antagonistic stiffness will act like a spring with a large spring constant. Upon disturbances, its response will have a lower magnitude. Also, a McKibben actuator at higher pressure can dissipate energy faster due to the larger internal friction between the fibers. Therefore, the joint can structurally lower the oscillation magnitude and increase the damping factor. As a result, the arm can quickly stabilize when a new target position is given.
IV-C Positional Control Precision
Figure 6: The UMArm is tasked with drawing circular and star-shaped patterns in the spatial coordinate system. The target trajectory is shown in red, with a point moving along it at constant speed. The commanded trajectory from the inverse kinematics controller is shown in green, and the actual end-effector trajectory recorded by the motion capture system is shown in blue.
Figure 7: (A) Experimental setup for testing UMArm’s force generation at different configurations. (B) The UMArm is first commanded to hold its end-effector at a fixed position. A Kinova Arm then applies push and pull motions to displace the end-effector from this position. In response, the UMArm exerts force on the push–pull pad as it attempts to return to the commanded position. The maximum generated force is measured by the force–torque sensor and plotted in each figure. This procedure is repeated for three different arm configurations.
We first evaluated the UMArm’s inverse kinematics controller in simulation, without connecting to the physical hardware. A total of 1000 points were randomly sampled within the UMArm workspace, and the UMArm was tasked with traversing these waypoints using its end-effector. A computation time of 0.5 s was allowed between consecutive waypoints. The resulting position errors showed strong accuracy: 50% of the points were within 0.1 mm, 91% within 1 mm, 99% within 2 mm, and all points within 3 mm.
Next, we conducted a trajectory tracking experiment with the UMArm hardware. As shown in Figure 6, the arm was tasked with drawing (circle and star) in the spatial coordinates. The desired trajectory was first generated, and then a target point was moved along the pattern at constant speed. The inverse kinematics controller then guided the UMArm’s end-effector to follow this moving target. The figure plots the reference pattern, the commanded trajectory from the controller, and the actual trajectory recorded by the motion capture system.
We tested three conditions for each pattern: (i) no load at standard speed, (ii) no load at a higher speed, and (iii) 500 g load at standard speed. The tracking error is defined as the Euclidean distance between the target position and the actual position measured by motion capture. Because the joint controller is PID-based, some error is unavoidable, primarily due to the lag of the actual position behind the moving target. Nonetheless, the tracking remains smooth and stable, and the overall patterns drawn are very close to the intended trajectories.
IV-D Force Generation Capability
The force generation capability of the UMArm is characterized experimentally. Figure 7A shows the setup. A force–torque sensor (ATI Nano-17) is mounted on the UMArm, with a spherical contact dome attached to the sensor. A Kinova Gen3 arm is then used to push and pull the contact dome. Three end-effector positions are tested, and at each position, the UMArm’s ability to apply force in both pushing and pulling directions is evaluated.
Figure 7B illustrates the experimental procedure. The UMArm first uses its inverse kinematics controller to move the end-effector to one of three predefined positions. The Kinova Arm then displaces the end-effector by 50 mm, either pushing or pulling it away from the target position. Owing to its controller design, the UMArm attempts to restore the end-effector position, thereby applying a reaction force normal to the push–pull pad of the Kinova Arm. This force is measured by the force–torque sensor.
At position 1, the arm generates similar force outputs in both pushing and pulling directions. At position 2, it produces a greater force in pulling and a reduced force in pushing, while at position 3 this asymmetry is even more pronounced. The difference arises from actuator length asymmetry: at position 1, all actuators are approximately equal in length, whereas at positions 2 and 3 the actuators responsible for pushing are shortened while those for pulling are stretched. Since McKibben actuators generate stronger forces when stretched at the same pneumatic pressure, larger pulling forces than pushing forces are observed at positions 2 and 3.
IV-E Untethered Operation
@@
The arm can operate untethered using a portable pressure source. We demonstrate this functionality using an off-the-shelf compressed air tank that holds 1.11L (68ci) at 31000kPa (4500 psi) (Fig 11A). It was originally built for airsoft players, so it is designed to be portable and lightweight. We characterize the untethered performance by letting the robot traverse the eight way-points with two 2-second intervals.The way-point arrangement is shown in Fig. 2 in the Supplementary Material. The experiment results show that the compressed air tank we use could supply the robot to traverse 700 waypoints, or run continuously for 11min 45s before it starts to lose tracking accuracy due to the depletion of the tank.
@@ V UMArm Demonstrations
V-A Variable Structural Stiffness
Figure 8: Testing UMArm’s variable structural stiffness ability. (A1, A2). The McKibben actuators are selectively pressurized for different tool-frame stiffness profiles. A Kinova Gen3 Arm is equipped with a frame with printer paper installed. A 2-direction spike is installed on the force-torque sensor on the end-effector of the UMArm. (B1, B2). The Kinova Arm pushes the paper frame against the 2-direction spike. The spike can easily penetrate the paper along the stiff direction, but cannot penetrate the paper along the soft direction. The force for each case is recorded by the force torque sensor. The skeleton model plots also show the estimated vs actual displacement from the applied external force.
Figure 9: A 2-DoF end-effector module consists of 2 encoder motors and a gripper is installed on the arm. Using the arm’s current configuration, the controller will calculate encoder motor angles so that the motors can maintain the spatial orientation of the gripper. The gripper is driven by a mini servo. B: With the installed gripper end-effector, the arm can stack the three objects together. In the fifth and sixth sub-figures, an extra object is placed ahead of time by a person for the next round of stacking.
@@ V-A Variable Structural Stiffness
Soft robotic arms are structurally compliant, which is crucial for safe interaction with humans. However, for a robotic arm to be precise and fast, it needs to have control over its structural stiffness. Otherwise, we might see oscillatory behavior or buckling under heavy load. UMArm can control the structural compliance on all 12 joints on the arm. This has granted the UMArm great flexibility to have safety and performance simultaneously. Figure 8 demonstrates UMArm’s directional compliance control capability.
We 3D printed a two-direction spike and mounted it perpendicularly on the end-effector. By selectively stiffening certain joints and relaxing others, the UMArm was configured to be structurally stiff along one direction while remaining soft along the other direction. To observe this anisotropic stiffness, a Kinova Arm held a sheet of printer paper and pressed it against the spike.
Figure 8 A1/B1 show the UMArm under the first stiffness profile, where the end-effector is stiff in the left–right direction and soft in the front–back direction. A2/B2 illustrate the inverse profile. In both cases, the printer paper was not penetrated when pressed against the soft direction, but was penetrated when pressed against the stiff direction.
To further analyze UMArm’s variable stiffness capability, we would like to develop a modeling approach to quantitatively estimate directional compliance as a function of the arm’s configuration and actuator pressures. The structure of UMArm is fully compatible with the stiffness estimation method proposed by Bruder et al. [bruder2023increasing]. We adapted the principles of their approach and implemented a three-dimensional version tailored to UMArm.
Using the configuration vector q and the actuator pressure vector p, we compute the task-space compliance matrix $C_{\text{x}}(\textbf{q},\textbf{p})\in\mathbb{R}^{3\times 3}$. The derivation of $C_{\text{x}}$ is provided in the Supplementary Materials. This matrix can be used to predict the virtual displacement $\delta\textbf{x}\in\mathbb{R}^{3}$ of the end-effector under an external virtual force $\delta F_{\text{ext}}\in\mathbb{R}^{3}$:
$\delta\textbf{x}=C_{\textbf{x}}(\textbf{q,p})\delta F_{\text{ext}}$ (4)
During the paper-penetration test, external force data were collected from the force–torque sensor, while displacement data were obtained from the motion capture system. These measurements were used to validate the variable stiffness model. In Figure 8 B1/B2, the skeleton model of UMArm prior to each paper press is shown in gray. For the soft direction, force and displacement are measured after the system stabilizes; for the stiff direction, they are measured just before the spikes penetrate the paper. Externally applied forces are plotted as orange dashed arrows. Although the skeleton model is shown in the pre-displacement state, the plotted force vectors were recorded after displacement. The post-displacement force vector is transformed in the spatial frame to preserve its relative alignment with the robot. The actual displacement vector is shown as a light blue arrow, starting from the pre-displacement end-effector position and ending at the post-displacement position. The task-space compliance matrix $C_{\text{x}}$ is computed using the pre-displacement configuration q; when multiplied by the plotted force vector, it yields the estimated displacement, shown as a medium blue arrow.
Although the estimated displacement is valid only locally, the results still reveal a clear trend: along the compliant direction, small forces produce relatively large displacements, whereas along the stiff direction, large forces result in only small displacements. These results confirm that the modeled structural compliance matrix provides a useful quantitative estimate of the UMArm’s end-effector stiffness profile.
V-B Precision Handling Arm